OpenStudy (anonymous):

1/2x+1/3y=3 Which of the following equations is equivalent to the given equation? 3x + 2y = 3 2x + 3y = 18 3x + 2y = 18

OpenStudy (anonymous):

3x + 2y = 3 2x + 3y = 18 3x + 2y = 18